The average height of male in the Philippines with age 20 to 39 is 163.5 cm while the standard deviation is 37.8 cm. If we choose 200 from these people what will be the standard deviation of this sample


1
Expert's answer
2021-07-02T07:20:18-0400

By using central limit theorem,

"\mu_{\bar{x}}=\mu" and

"\sigma_{\bar{x}}=\frac{\sigma}{\sqrt{n}}"

Given

"\mu=163.5, \sigma=37.8, n=200" , then

"\mu_{\bar{x}}=163.5"

"\sigma_{\bar{x}}=\frac{37.8}{\sqrt{200}}=2.67"

The standard deviation will be 2.67